OpenStudy (anonymous):

Tap A can fill a cylindrical container in 3 minutes. Tap B can fill the same container in 15 minutes. How long would it take the two taps together to fill the container? I always forget how to do these type of questions, please provide the solution, not only the result :) Thanks!

OpenStudy (anonymous):

I got it tap A fils in one minute = 1/3 cylinder tap B fills in one minute = 1/15 cylinder both in one minute = 1/3 + 1/15 = 2/5 cylinder both fill full cylinder in = 5/2 minute = 2 minute and 20 seconds